I am trying to figure out how to write a mathematical statement for the following question:
"Jill was playing a game with her sister. Jill was down by 5. After her next turn, she was down by 3. Write a mathematical statement that represents this scenario."
Should the statement be written as x - 5, x - 3 ?

Answer by Theo(13342)

y = -2x + 7 for x = 1 to 2

y represents the number of points she is down by.
x represents the number of turns.

when x = 1, y = 7 - 2 = 5
when x = 2, y = 7 - 4 = 3